# Using quantum computers to identify prime numbers via entanglement dynamics

### v1

@@ Abstract
Normally, Eq. (19) would be an integral over the whole period $T$, but we are employing the property of the symmetry of $\gamma_{A}(t)$, presented earlier in this letter. After calculating the Fourier modes $\alpha_{n}$, the last part of our algorithm involves comparing the value of $\alpha_{n}$ with the analytical lower bound $B_{n}$. In this final step, the numerical integration is done in $p$ partitions, resulting in an equivalent number of points used for $\gamma_{A}(t)$ in the interval $20\leq t\leq T/$ 2. Consequently, to achieve a desired precision $\epsilon$, our quantum circuit requires at least $p$ executions. Currently, the exact scaling of $p$ with respect to $d$, for a given $\epsilon$, remains undetermined.
@@
Simulations.— To evaluate the applicability of our algorithm, simulations were performed using IBM’s Qiskit framework (version 0.45.1). These simulations targeted three distinct values of $d$, with results depicted in blue in Fig. 2. For all the simulations, we used $10^{5}$ shots and fixed $\omega=0.1\,\text{s}^{-1}$, as changing the value of $\omega$ has no effect on the Fourier modes $\alpha_{n}$. Regarding the number of executions $p$ of the circuit, we selected $p=375$, $p=1500$, and $p=6000$ for the dimensions $d=16$, $d=32$, and $d=64$, respectively. The values of $p$ were chosen to achieve roughly the same accuracy for the three values of $d$. Using Python (version 3.11.3) with the Scipy library (version 1.11.3), Fourier modes $\alpha_{n}$ were calculated with Simpson’s rule for the numerical integration of Eq. (19). Due to the substantial size of the quantum circuit for the three dimensions analyzed in our simulations, we present the circuit for a lower dimension, $d=4$, purely for illustrative purposes. This simplified example is shown in Fig. 1, allowing us to convey the structure without the complexity of the larger dimensions.
Figure 2: Comparison of simulation results with theoretical predictions for the Fourier modes of the reduced purity across different dimensions $(d=16,32,64)$. Red points represent the analytical values of $\alpha_{n}$, calculated directly from their theoretical expressions, while the green line stands for the minimum value for the Fourier modes, also derived from theoretical calculations. Blue points illustrate the Fourier modes obtained through numerical integration of the reduced purity $\gamma_{A}(t)$ extracted from the classical emulation of our quantum circuit. The numerical integration was performed using various partition values $(p=375,1500,6000)$. Prime numbers are expected to align with the lower bound $B_{n}$, whereas composite numbers appear above. Points closely approaching the lower bound $B_{n}$, yet not precisely on it, may be considered prime candidates, subject to verification due to the presence of statistical error.
Conclusions.— This work presents a qubit-based quantum algorithm for prime number identification, rooted in the analysis of entangled subsystem dynamics. By strategically selecting a bipartite Hamiltonian of dimension $d^{2}$ characterized by energy quantization, we demonstrate that the Fourier modes $\alpha_{n}$ of the reduced purity display distinctive patterns. These patterns enable the differentiation between prime and composite numbers for any integer $n$ such that $1\leq n\leq 2(d-1)$. To implement this protocol as an algorithm for qubit-based quantum computing, we map this $2$-qudit system into a $q$-qubit system using the expression $d^{2}=2^{q}$. In this mapping, as an independent result, we found that the local unitary gates with the form of Eq. (2) may be implemented in polynomial time using the method of Ref. welch .
Executing our quantum circuit involves three stages, as illustrated in Fig. 1: preparation of the initial state $|\psi(0)\rangle_{AB}$ for both sets of $q$ qubits, preparation of the evolved state $|\psi(t)\rangle_{AB}$ also for both sets, and estimating the reduced purity $\gamma_{A}(t)$. These stages have respective gate costs: $2G_{1}=2q$, $2G_{2}(q)=\frac{3}{2}q^{2}+q$, and $G_{3}=\frac{3}{2}q+2$. In particular, $G_{2}(q)$ represents the number of gates necessary to implement local unitary gates that share the same mathematical structure as $\hat{U}(t)$, where this implementation is accomplished using the methodology outlined in Ref. welch . Contrary to the anticipated exponential cost of the method delineated in that work, we have demonstrated that $G_{2}(q)$ scales polynomially with the number of qubits $q$, thereby affirming the high efficiency of that method in preparing the class of states $|\psi(t)\rangle_{AB}$, as defined by Eq. (3). Thus, the total gate cost $G(q)$ for one execution of our quantum circuit is given by $G(q)=\frac{3}{2}q^{2}+\frac{11}{2}q+2$. By defining $N=2(d-1)$ and using $2^{q}=d^{2}$, we find that the asymptotic gate cost is $\mathcal{O}(\log^{2}_{10}(N))$, indicating a quadratic scaling in the number of digits of $N$.
Naturally, although this work presented ideal classical simulations, our algorithm could be implemented on quantum hardware. One major challenge that may arise, however, is that controlled-NOT and SWAP gates require good connectivity between the qubits, which is not always available in existing superconducting qubit-based quantum computers. A simple solution is to explore alternative quantum computing technologies. Trapped ion quantum computers, exemplified by those developed by IonQ and Quantinuum, offer extended coherence times and all-to-all connectivity between qubits, allowing for more effective implementation of these gates. Another approach would be to individually modify the last two stages. For the second stage, which involves preparing $|\psi(t)\rangle_{AB}$ for both copies, alternative methods could be considered, although we suspect that Ref. welch provides the most efficient technique for this purpose. For the third and final stage, since the reduced purity $\gamma_{A}(t)$ is related to the reduced density operator $\hat{\rho}_{A}(t)$, we could replace the direct measurement of $\gamma_{A}(t)$ with a state tomography of $\hat{\rho}_{A}(t)$. Modern tomographic methods, such as classical shadows tomography Aaronson2018 ; Huang2020 ; Nguyen2022 ; Bu2024 , show promise and can be employed for the estimation of $\gamma_{A}(t)$, requiring only $q$ qubits instead of $12q+$ 1.
Importantly, our demonstration that the unitary gate $\hat{U}(t)$ can be implemented with a polynomial number of gates, contrary to the general expectation of exponential gate requirements, paves the way for the efficient realization of similar transformations in quantum computing. This finding not only enhances the feasibility of executing our prime number identification algorithm on actual quantum systems but also suggests a broader applicability for optimizing other quantum algorithms. As quantum computing continues to evolve, the adaptation of such efficient unitary operations will be crucial in overcoming the limitations posed by gate complexities and in harnessing the full potential of quantum technologies. With advancements in fault tolerance and increased access to a larger number of qubits in the future, our algorithm holds the potential to verify the primality of significantly larger primes, marking a significant step forward in the practical application of quantum algorithms for fundamental computational problems.

### v2

@@ Abstract
Normally, Eq. (17) would be an integral over the whole period $T$, but we are employing the property of the symmetry of $\gamma_{A}(t)$, presented earlier in this article. After calculating the Fourier modes $\alpha_{n}$, the last part of our algorithm involves comparing the value of $\alpha_{n}$ with the analytical lower bound $B_{n}$. In this final step, the numerical integration is done in $p$ partitions, resulting in an equivalent number of points used for $\gamma_{A}(t)$ in the interval $20\leq t\leq T/$ 2. Consequently, to achieve a desired precision $\epsilon$, our quantum circuit requires at least $p$ executions. Currently, the exact optimal scaling of $p$ with respect to $d$, for a given $\epsilon$, remains undetermined.
Figure 1: Comparison of simulation results with theoretical predictions for the Fourier modes of the reduced purity across different dimensions $(d=16,32,64)$. Red points represent the analytical values of $\alpha_{n}$, calculated directly from their theoretical expressions, while the green line stands for the minimum value for the Fourier modes, also derived from theoretical calculations. Blue points illustrate the Fourier modes obtained through numerical integration of the reduced purity $\gamma_{A}(t)$ extracted from the classical emulation of our quantum circuit. The numerical integration was performed using various partition values $(p=375,1500,6000)$. Prime numbers are expected to align with the lower bound $B_{n}$, whereas composite numbers appear above.
In order to evaluate the applicability of our algorithm, classical simulations were performed using IBM’s Qiskit framework (version 0.45.1). These simulations targeted three distinct values of $d$, with results depicted in blue in Fig. 1. For all the simulations, we used $10^{5}$ shots and fixed $\omega=0.1\,\text{s}^{-1}$, as changing the value of $\omega$ has no effect on the Fourier modes $\alpha_{n}$. Regarding the number of executions $p$ of the circuit, we selected $p=375$, $p=1500$, and $p=6000$ for the dimensions $d=16$, $d=32$, and $d=64$, respectively. The values of $p$ were chosen to achieve roughly the same accuracy for the three values of $d$. Using Python (version 3.11.3) with the Scipy library (version 1.11.3), Fourier modes $\alpha_{n}$ were calculated with Simpson’s rule for the numerical integration of Eq. (17). Due to the substantial size of the quantum circuit for the three dimensions analyzed in our simulations, we present the circuit for a lower dimension, $d=4$, purely for illustrative purposes. This simplified example is shown in Fig. 2, allowing us to convey the structure without the complexity of the larger dimensions.
@@
Concluding, this work presented a qubit-based quantum algorithm for prime number identification, rooted in the analysis of entangled subsystem dynamics. By employing a bipartite Hamiltonian and analyzing the Fourier modes of the reduced purity, we distinguish between prime and composite numbers within the range $2\leq n\leq 2(d-1)$. Implementing this on a qubit-based system involves transforming a 2-qudit system into a qubit system, with the local unitary gate of Eq. (2) implementable in polynomial time, contrary to the expected exponential gate requirements.
Our quantum circuit executes in three stages with a total gate cost indicating quadratic scaling in the number of digits of $N=2(d-1)$. Despite idealized simulations, implementation on quantum hardware is feasible but faces challenges such as qubit connectivity. Alternatives like trapped ion quantum computers or modified gate preparation and reduced purity measurement methods could overcome these.
The efficient realization of unitary operations demonstrates the potential for broader application in quantum computing, suggesting future work could extend this algorithm to verify larger primes. This progress in quantum algorithm optimization could significantly impact the field’s practical application to fundamental computational problems.